OpenStudy (anonymous):

can someone help me please Write 243 in exponential form

OpenStudy (anonymous):

\[ \huge e^ {\ln 243 } \]

OpenStudy (lalaly):

\[3^5\]

OpenStudy (anonymous):

Haha. I think she means like 3^5.

OpenStudy (anonymous):

FSM, Do you mean what I posted is not an exponential form? :P

OpenStudy (anonymous):

Haha but doesn't \[\Large {e^{\ln 243}} = {243^{{e^{\ln 1}}}} \cdot \tan 45\]

OpenStudy (anonymous):

can anyone help me on this